How many gallons of a 9% peroxide solution must be mixed with 3 gallons of a 15% peroxide solution to obtain a 12% solution?
Answer by stanbon(75887)

How many gallons of a 9% peroxide solution must be mixed with 3 gallons of a 15% peroxide solution to obtain a 12% solution?
--------------------
Equation:
active + active = active
0.09x + 0.15*3 = 0.12(x+3)
------
Multiply thru by 100 to get:
9x + 15*3 = 12x + 12*3
-3x = -3*3
---
x = 3 gallons (amt of 9% solution needed in the mixture)
==============================
Cheers,
Stan H.